The market for a chemical pesticide can be modeled as given below, where Q is measured in thousands of pounds:

            MSB = 20 – 0.4Q                    MPB = 20 – 0.3Q                    MSC = MPC = 12 + 0.1Q

            Use this model to respond to the following questions.

a.         Find the efficient output level for this market.

b.         Find the equation that represents the negative consumption externality in this case.

c.         Determine the value of a product charge that would achieve the efficient solution.
Answer #1

A)

Efficient production takes place at the point: MSB = MSC

That is, 20-0.4Q = 12+0.1Q

That is, 8 = 0.5Q

Q* = 16

B)

Negative externality = MSB-MPB = 20-0.4Q - (20-0.3Q) = -0.1Q

This is the negative externality

C)

Value of product to be charged to achieve efficient allocation:

P* = 20-0.4(16) = 12+0.1(16) = $13.6